OpenStudy (anonymous):

in triangle ABC, a=600m, angle A=43.82degrees, and angle C=57.64degrees. find the length of the side opposite angle B.

OpenStudy (anonymous):

B=180-43.82-57.64= 78.54 degree a/sinA=b/sinB 600/sin43.82=b/sin78.54 you calculate b yourself.